# Determination of the stably free cancellation property for orders

### v2

@@ 5.5. A criterion assuming the Eichler condition for one component
Let $\mathfrak{I}$ be a two-sided ideal of $\Lambda$ such that $\Lambda/\mathfrak{I}$ is $R$-torsion and let $f:\Lambda\rightarrow\Lambda/\mathfrak{I}$ be the canonical quotient map. Assume that $A$ satisfies the Eichler condition. Then $f(\Lambda^{\times})$ contains the commutator subgroup $[(\Lambda/\mathfrak{I})^{\times},(\Lambda/\mathfrak{I})^{\times}]$.
@@
$[\overline{\Lambda}^{\times},\overline{\Lambda}^{\times}]\trianglelefteq g_{1}(\Lambda_{1}^{\times})\trianglelefteq W\trianglelefteq\ker(\partial)\trianglelefteq\overline{\Lambda}^{\times},$
@@
The hypotheses imply that we are in the setting of Example 5.1. In particular, $\overline{\Lambda}$ is $R$-torsion and hence finite. Moreover, every subgroup of $\overline{\Lambda}^{\times}$ containing $[\overline{\Lambda}^{\times},\overline{\Lambda}^{\times}]$ is necessarily normal in $\overline{\Lambda}^{\times}$ and has finite abelian quotient. Theorem 5.10 applied with $f=g_{1}$ implies that $[\overline{\Lambda}^{\times},\overline{\Lambda}^{\times}]\trianglelefteq g_{1}(\Lambda_{1}^{\times})$. Since $W=g_{1}(\Lambda_{1}^{\times})g_{2}(\Lambda_{2}^{\times})$ by definition and $g_{1}(\Lambda_{1}^{\times})$ is normal in $\overline{\Lambda}^{\times}$, we have that $W$ is in fact a subgroup of $\overline{\Lambda}^{\times}$ (a priori, it is only a subset), and that $g_{1}(\Lambda_{1}^{\times})\trianglelefteq W$. The remaining containments follow from Corollary 5.3. ∎
@@
Assume the setting of Proposition 5.11. Then $W$ is a group. Suppose that $N$ is a subgroup of $W$ containing $[\overline{\Lambda}^{\times},\overline{\Lambda}^{\times}]$. Then $N\trianglelefteq\ker(\partial)$ and $\ker(\partial)/N$ is a finite abelian group. Let $S$ be a generating subset of $\ker(\partial)/N$. Suppose that $\Lambda_{2}$ has SFC. Then the following assertions are equivalent:
@@ 6. Applications to integral group rings
6.1. A fibre product for integral group rings
We first record the following well-known instance of a fibre product. Let $G$ be a finite group, let $N$ be a normal subgroup and let $H=G/N$. Let $n=|N|$ and let $\mathrm{Tr}_{N}=\sum_{\tau\in N}\tau\in\mathbb{Z}[G]$ be the trace element associated to $N$. Let $\pi_{1}:\mathbb{Z}[G]\rightarrow\mathbb{Z}[H]$ be the canonical projection and let $I=\ker(\pi_{1})$. Let $J=\mathrm{Tr}_{N}\mathbb{Z}[G]$, let $\Gamma=\mathbb{Z}[G]/J$ and let $\pi_{2}:\mathbb{Z}[G]\rightarrow\mathbb{Z}[G]/J$ be the canonical projection. Then $I\cap J=0$ and $I+J=n\mathbb{Z}[G]$. Thus, by Example 5.1 we have a fibre product
${\mathbb{Z}[G]}$${\mathbb{Z}[H]}$${\Gamma}$${(\mathbb{Z}/n\mathbb{Z})[H],}$$\scriptstyle{\pi_{2}}$$\scriptstyle{\pi_{1}}$$\scriptstyle{g_{1}}$$\scriptstyle{g_{2}}$ (6.1)
@@ 6.1. A fibre product for integral group rings
$|W_{G,N}|\leq\frac{|\mathrm{Cl}(\mathbb{Z}[H])||\mathrm{Cl}(\mathbb{Z}[\Gamma])||(\mathbb{Z}/n\mathbb{Z})[H]^{\times}|}{|\mathrm{Cl}(\mathbb{Z}[G])|}.$ (6.2)
@@ 6.2. Integral group rings of the form $\mathbb{Z}[H\times C_{2}]$
    $|\mathbb{F}_{2}[Q_{8}]^{\times}|=2^{7}$ (since $\mathbb{F}_{2}[Q_{8}]$ is a local ring).
@@
and so $\mathbb{Z}[Q_{8}\times C_{2}]$ does not have SFC. Note that the quantities in (iii) and (iv) can instead both be calculated on a computer by using the algorithm of [BB06, §3].
@@
The algorithm of [BB06, §3] shows that $|\mathrm{Cl}(\mathbb{Z}[\tilde{T}])|=2$ and $|\mathrm{Cl}(\mathbb{Z}[\tilde{T}\times C_{2}])|=2^{6}$. By Remark 5.9, we have $|\mathbb{F}_{2}[\tilde{T}]^{\times}|=|J(\mathbb{F}_{2}[\tilde{T}])||(\mathbb{F}_{2}[\tilde{T}]/J(\mathbb{F}_{2}[\tilde{T}]))^{\times}|$, where $J(\mathbb{F}_{2}[\tilde{T}])$ denotes the Jacobson radical of $\mathbb{F}_{2}[\tilde{T}]$. By [MS22, Theorem 2.5] we have $\mathbb{F}_{2}[\tilde{T}]/J(\mathbb{F}_{2}[\tilde{T}])\cong\mathbb{F}_{2}\oplus\mathbb{F}_{4}$ (this can also be verified using the algorithm of [FR85]). Hence $|J(\mathbb{F}_{2}[\tilde{T}])|=2^{24-3}=2^{21}$. Moreover, $(\mathbb{F}_{2}[\tilde{T}]/J(\mathbb{F}_{2}[\tilde{T}]))^{\times}\cong(\mathbb{F}_{2}\oplus\mathbb{F}_{4})^{\times}\cong C_{3}$. Therefore $|\mathbb{F}_{2}[\tilde{T}]^{\times}|=3\cdot 2^{21}$.
The integral group ring $\mathbb{Z}[\tilde{T}]$ has SFC by [Swa83, Theorem I]. Let $V=V_{\tilde{T}}$ be the image of the canonical projection $\pi:\mathbb{Z}[\tilde{T}]^{\times}\rightarrow\mathbb{F}_{2}[\tilde{T}]^{\times}$. Then by Proposition 6.4 and the above calculations we have
@@
Then $u_{g,h}$ and $u^{\prime}_{g,h}$ are elements of $\mathbb{Z}[\tilde{T}]^{\times}$. Let
$U=\langle g,\pi(u_{g,h}),\pi(u^{\prime}_{g,h})\mid g,h\in\tilde{T}\rangle$
and note that this is a subgroup of $V$. By Remark 6.3, we obtain an explicit embedding of $\mathbb{F}_{2}[\tilde{T}]^{\times}$ into $\mathrm{GL}_{24}(\mathbb{F}_{2})$. Since a calculation on a computer algebra system such as Magma [BCP97] or Oscar [OSC24, DEF+24, FHHJ17] shows that the image of $U$ under this embedding is a subgroup of $\mathrm{GL}_{24}(\mathbb{F}_{2})$ of order $3\cdot 2^{17}$, we must have $U=V$ and $|V|=3\cdot 2^{17}$. Therefore we conclude that $\mathbb{Z}[\tilde{T}\times C_{2}]$ has SFC. ∎
@@ 6.3. A useful reduction criterion for certain integral group rings
The following reduction criterion only uses the sizes of certain locally free class groups, and thus can be applied relatively easily. It will be used in the proof of Theorem 13.1 (i.e. Theorem 1.3).
@@
Let $G$ be a finite group with a normal subgroup $N$ of size $2$ and let $H=G/N$. Let $\Gamma=\mathbb{Z}[G]/\mathrm{Tr}_{N}\mathbb{Z}[G]$. Suppose that $\mathbb{Z}[H\times C_{2}]$ has SFC and that
$\frac{|\mathrm{Cl}(\mathbb{Z}[H\times C_{2}])|}{|\mathrm{Cl}(\mathbb{Z}[H])|}=\frac{|\mathrm{Cl}(\mathbb{Z}[G])|}{|\mathrm{Cl}(\Gamma)|}.$ (6.3)
@@
$|V_{H}|=\frac{|\mathrm{Cl}(\mathbb{Z}[H])|^{2}|\mathbb{F}_{2}[H]^{\times}|}{|\mathrm{Cl}(\mathbb{Z}[H\times C_{2}])|}=\frac{|\mathrm{Cl}(\mathbb{Z}[H])||\mathbb{F}_{2}[H]^{\times}||\mathrm{Cl}(\Gamma)|}{|\mathrm{Cl}(\mathbb{Z}[G])|}.$ (6.4)
@@ 7.1. Setup
Let $K$ be a number field with ring of integers $\mathcal{O}=\mathcal{O}_{K}$, and let $A$ be a finite-dimensional semisimple $K$-algebra. Let $\Lambda$ be an $\mathcal{O}$-order in $A$. By [Rei03, (10.4)] there exists a (not necessarily unique) maximal $\mathcal{O}$-order $\mathcal{M}$ in $A$ containing $\Lambda$. Let $\mathfrak{f}$ be any proper full two-sided ideal of $\mathcal{M}$ that is contained in $\Lambda$.
@@ 7.2. A criterion for freeness
$\pi:(\mathcal{M}/\mathfrak{f})^{\times}\longrightarrow{(\mathcal{M}/\mathfrak{f})^{\times}}/{(\Lambda/\mathfrak{f})^{\times}}$ (7.1)
@@
[BHJ22, Proposition 4.3 and Theorem 4.5] only assumed that $\beta\in\mathcal{M}$ but should have assumed that $\beta\in\mathcal{M}\cap A^{\times}$, or equivalently, that $\beta\in\mathcal{M}$ and that $X$ is a full left ideal of $\Lambda$, that is, $KX=A$. Note that any locally free left ideal of $\Lambda$ is necessarily full. The strengthened assumption is needed to ensure that the condition $X=\Lambda\alpha$ implies that $X$ is in fact free over $\Lambda$ (see [BHJ22, Lemma 2.3]); without this, the implication (i) $\Rightarrow$ (v) in Proposition 7.1 may not hold. This omission does not affect any of the other results of [BHJ22].
@@ 7.3. An exact sequence
Most of the rest of this section shall be concerned with the development of a criterion for stable freeness. We shall first need a certain exact sequence.
@@
Let $g_{1}=g_{2}:\Lambda\rightarrow\Lambda/\mathfrak{f}$ be the canonical projection map. For $i=1,2$ define $\pi_{i}:\mathcal{D}\rightarrow\Lambda$ to be the restriction of the projection map from $\Lambda\oplus\Lambda$ onto its $i$th component. Then we have a fibre-product
${\mathcal{D}}$${\Lambda}$${\Lambda}$${\Lambda/\mathfrak{f}.}$$\scriptstyle{\pi_{2}}$$\scriptstyle{\pi_{1}}$$\scriptstyle{g_{1}}$$\scriptstyle{g_{2}}$ (7.2)
For $i=1,2$ define $q_{i}:\mathrm{Cl}(\mathcal{D})\rightarrow\mathrm{Cl}(\Lambda)$ by $[X]_{\mathrm{st}}\mapsto[\Lambda\otimes_{\mathcal{D},\pi_{i}}X]_{\mathrm{st}}$. Then by Proposition 5.6 we obtain an exact sequence
@@
$(\Lambda/\mathfrak{f})^{\times}\stackrel{{\scriptstyle\partial}}{{\longrightarrow}}\mathrm{Cl}(\Lambda,\mathfrak{f})\stackrel{{\scriptstyle q_{1}}}{{\longrightarrow}}\mathrm{Cl}(\Lambda)\longrightarrow 0.$ (7.3)
@@ 7.4. Decomposition via central idempotents
Much of the following notation is adopted from [BHJ22]. Denote the centre of a ring $R$ by $Z(R)$. Let $C=Z(A)$ and let $\mathcal{O}_{C}$ be the integral closure of $\mathcal{O}$ in $C$. Set $\mathfrak{g}:=\mathfrak{f}\cap\mathcal{O}_{C}$ and note that this is a proper full ideal of $\mathcal{O}_{C}$. Let $e_{1},\ldots,e_{r}$ be the primitive idempotents of $C$ and set $A_{i}=Ae_{i}$. Then
$A=A_{1}\oplus\cdots\oplus A_{r}$ (7.4)
is a decomposition of $A$ into indecomposable two-sided ideals (see [CR81, (3.22)]). Each $A_{i}$ is a simple $K$-algebra with identity element $e_{i}$. The centres $K_{i}:=Z(A_{i})$ are finite field extensions of $K$ via $K\rightarrow K_{i}$, $\alpha\mapsto\alpha e_{i}$, and we have $K$-algebra isomorphisms $A_{i}\cong\mathrm{Mat}_{n_{i}}(D_{i})$ where $D_{i}$ is a division algebra with $Z(D_{i})\cong K_{i}$ (see [CR81, (3.28)]). The Wedderburn decomposition (7.4) induces decompositions
$C=K_{1}\oplus\cdots\oplus K_{r}\quad\textrm{and}\quad\mathcal{O}_{C}=\mathcal{O}_{K_{1}}\oplus\cdots\oplus\mathcal{O}_{K_{r}},$ (7.5)
@@
$\mathcal{M}=\mathcal{M}_{1}\oplus\cdots\oplus\mathcal{M}_{r},\quad\mathfrak{f}=\mathfrak{f}_{1}\oplus\cdots\oplus\mathfrak{f}_{r}\quad\textrm{and}\quad\mathfrak{g}=\mathfrak{g}_{1}\oplus\cdots\oplus\mathfrak{g}_{r}.$ (7.6)

### v3

@@ 5.5. A criterion assuming the Eichler condition for one component
Let $\mathfrak{I}$ be a two-sided ideal of $\Lambda$ such that $\Lambda/\mathfrak{I}$ is $R$-torsion and let $f:\Lambda\rightarrow\Lambda/\mathfrak{I}$ be the canonical quotient map. Assume that $A$ satisfies the Eichler condition. Then $f(\Lambda^{\times})$ contains the commutator subgroup $(\Lambda/\mathfrak{I})^{\times},(\Lambda/\mathfrak{I})^{\times}]$.
@@
$\overline{\Lambda}^{\times},\overline{\Lambda}^{\times}]\trianglelefteq g_{1}(\Lambda_{1}^{\times})\trianglelefteq W\trianglelefteq\ker(\partial)\trianglelefteq\overline{\Lambda}^{\times},$
@@
The hypotheses imply that we are in the setting of Example 5.1. In particular, $\overline{\Lambda}$ is $R$-torsion and hence finite. Moreover, every subgroup of $\overline{\Lambda}^{\times}$ containing $\overline{\Lambda}^{\times},\overline{\Lambda}^{\times}]$ is necessarily normal in $\overline{\Lambda}^{\times}$ and has finite abelian quotient. Theorem 5.10 applied with $f=g_{1}$ implies that $\overline{\Lambda}^{\times},\overline{\Lambda}^{\times}]\trianglelefteq g_{1}(\Lambda_{1}^{\times})$. Since $W=g_{1}(\Lambda_{1}^{\times})g_{2}(\Lambda_{2}^{\times})$ by definition and $g_{1}(\Lambda_{1}^{\times})$ is normal in $\overline{\Lambda}^{\times}$, we have that $W$ is in fact a subgroup of $\overline{\Lambda}^{\times}$ (a priori, it is only a subset), and that $g_{1}(\Lambda_{1}^{\times})\trianglelefteq W$. The remaining containments follow from Corollary 5.3. ∎
@@
Assume the setting of Proposition 5.11. Then $W$ is a group. Suppose that $N$ is a subgroup of $W$ containing $\overline{\Lambda}^{\times},\overline{\Lambda}^{\times}]$. Then $N\trianglelefteq\ker(\partial)$ and $\ker(\partial)/N$ is a finite abelian group. Let $S$ be a generating subset of $\ker(\partial)/N$. Suppose that $\Lambda_{2}$ has SFC. Then the following assertions are equivalent:
@@ 6. Applications to integral group rings
6.1. A fiber product for integral group rings
We first record the following well-known instance of a fiber product. Let $G$ be a finite group, let $N$ be a normal subgroup and let $H=G/N$. Let $n=|N|$ and let $\mathrm{Tr}_{N}=\sum_{\tau\in N}\tau\in\mathbb{Z}[G]$ be the trace element associated to $N$. Let $\pi_{1}:\mathbb{Z}[G]\rightarrow\mathbb{Z}[H]$ be the canonical projection and let $I=\ker(\pi_{1})$. Let $J=\mathrm{Tr}_{N}\mathbb{Z}[G]$, let $\Gamma=\mathbb{Z}[G]/J$ and let $\pi_{2}:\mathbb{Z}[G]\rightarrow\mathbb{Z}[G]/J$ be the canonical projection. Then $I\cap J=0$ and $I+J=n\mathbb{Z}[G]$. Thus, by Example 5.1 we have a fiber product
$\mathbb{Z}[G]$$\mathbb{Z}[H]$$\Gamma$$(\mathbb{Z}/n\mathbb{Z})[H],$$\scriptstyle{\pi_{2}}$$\scriptstyle{\pi_{1}}$$\scriptstyle{g_{1}}$$\scriptstyle{g_{2}}$
@@ 6.1. A fiber product for integral group rings
(6.1) $|W_{G,N}|\leq\frac{|\mathrm{Cl}(\mathbb{Z}[H])||\mathrm{Cl}(\mathbb{Z}[\Gamma])||(\mathbb{Z}/n\mathbb{Z})[H]^{\times}|}{|\mathrm{Cl}(\mathbb{Z}[G])|}.$
@@ 6.2. Integral group rings of the form $\mathbb{Z}[H\times C_{2}]$
    $|\mathbb{F}_{2}[Q_{8}]^{\times}|=2^{7}$ (since $\mathbb{F}_{2}[Q_{8}]$ is a local ring with residue field $\mathbb{F}_{2}$).
@@
and so $\mathbb{Z}[Q_{8}\times C_{2}]$ does not have SFC. Note that the quantities in (iii) and (iv) can instead be calculated on a computer by using the algorithm of [BB06, §3].
@@
[…]
The integral group ring $\mathbb{Z}[\tilde{T}$] has SFC by [Swa83, Theorem I]. Let $V=V_{\tilde{T}}$ be the image of the canonical projection $\pi:\mathbb{Z}[\tilde{T}]^{\times}\rightarrow\mathbb{F}_{2}[\tilde{T}$$^{\times}$. Then by Proposition 6.4 and the above calculations we have
@@
Then $g$, $u_{g,h}$ and $u^{\prime}_{g,h}$ are elements of $\mathbb{Z}[\tilde{T}$$^{\times}$. Let
$U=\langle\pi(g),\pi(u_{g,h}),\pi(u^{\prime}_{g,h})\mid g,h\in\tilde{T}\rangle$
and note that this is a subgroup of $V$. By Remark 6.3, we obtain an explicit embedding of $\mathbb{F}_{2}[\tilde{T}$$^{\times}$ into $\mathrm{GL}_{24}(\mathbb{F}_{2})$. Since a calculation on a computer algebra system such as Magma [BCP97] or Oscar [OSC25, DEF+25, FHHJ17] shows that the image of $U$ under this embedding is a subgroup of $\mathrm{GL}_{24}(\mathbb{F}_{2})$ of order $3\cdot 2^{17}$, we must have $U=V$ and $|V|=3\cdot 2^{17}$. Therefore we conclude that $\mathbb{Z}[\tilde{T}\times C_{2}]$ has SFC. ∎
@@ 6.3. A useful reduction criterion for certain integral group rings
The following reduction criterion only uses the orders of certain locally free class groups, and thus can be applied relatively easily. It will be used in the proof of Theorem 13.1 (i.e. Theorem 1.4).
@@
Let $G$ be a finite group with a normal subgroup $N$ of order $2$ and let $H=G/N$. Let $\Gamma=\mathbb{Z}[G]/\mathrm{Tr}_{N}\mathbb{Z}[G]$. Suppose that $\mathbb{Z}[H\times C_{2}]$ has SFC and that
(6.2) $\frac{|\mathrm{Cl}(\mathbb{Z}[H\times C_{2}])|}{|\mathrm{Cl}(\mathbb{Z}[H])|}=\frac{|\mathrm{Cl}(\mathbb{Z}[G])|}{|\mathrm{Cl}(\Gamma)|}.$
@@
(6.3) $|V_{H}|=\frac{|\mathrm{Cl}(\mathbb{Z}[H])|^{2}|\mathbb{F}_{2}[H]^{\times}|}{|\mathrm{Cl}(\mathbb{Z}[H\times C_{2}])|}=\frac{|\mathrm{Cl}(\mathbb{Z}[H])||\mathbb{F}_{2}[H]^{\times}||\mathrm{Cl}(\Gamma)|}{|\mathrm{Cl}(\mathbb{Z}[G])|}.$
@@ 7.1. Setup
Let $K$ be a number field with ring of integers $\mathcal{O}=\mathcal{O}_{K}$, and let $A$ be a finite-dimensional semisimple $K$-algebra. Let $\Lambda$ be an $\mathcal{O}$-order in $A$. By [Rei03, (10.4)] there exists a (not necessarily unique) maximal $\mathcal{O}$-order $\mathcal{M}$ in $A$ containing $\Lambda$. Let $\mathfrak{f}$ be any full two-sided ideal of $\mathcal{M}$ that is contained in $\Lambda$.
@@ 7.2. A criterion for freeness
(7.1) $\pi:(\mathcal{M}/\mathfrak{f})^{\times}\longrightarrow{(\mathcal{M}/\mathfrak{f})^{\times}}/{(\Lambda/\mathfrak{f})^{\times}}$
@@
[BHJ22, Proposition 4.3 and Theorem 4.5] only assumed that $\beta\in\mathcal{M}$ but should have assumed that $\beta\in\mathcal{M}\cap A^{\times}$, or equivalently, that $\beta\in\mathcal{M}$ and that $X$ is a full left ideal of $\Lambda$, that is, $KX=A$. Note that any locally free left ideal of $\Lambda$ is necessarily full. The strengthened assumption is needed to ensure that the condition $X=\Lambda\alpha$ implies that $X$ is in fact free over $\Lambda$ (see [BHJ22, Lemma 2.3]); without this, the implication (i) $\Rightarrow$ (v) in Proposition 7.1 may not hold. This omission does not affect any of the other results of [BHJ22]. Also note that in loc. cit. it was assumed that $\mathfrak{f}$ is strictly contained in $\mathcal{M}$, but in fact this assumption is unnecessary.
@@ 7.3. An exact sequence
Most of the rest of this section will be concerned with the development of a criterion for stable freeness. We shall first need a certain exact sequence.
@@
Let $g_{1}=g_{2}:\Lambda\rightarrow\Lambda/\mathfrak{f}$ be the canonical projection map. For $i=1,2$ define $\pi_{i}:\mathcal{D}\rightarrow\Lambda$ to be the restriction of the projection map from $\Lambda\oplus\Lambda$ onto its $i$th component. Then we have a fiber product
$\mathcal{D}$$\Lambda$$\Lambda$$\Lambda/\mathfrak{f}.$$\scriptstyle{\pi_{2}}$$\scriptstyle{\pi_{1}}$$\scriptstyle{g_{1}}$$\scriptstyle{g_{2}}$
For $i=1,2$ define $q_{i}:\mathrm{Cl}(\mathcal{D})\rightarrow\mathrm{Cl}(\Lambda)$ by $X]_{\mathrm{st}}\mapsto[\Lambda\otimes_{\mathcal{D},\pi_{i}}X]_{\mathrm{st}}$. Then by Proposition 5.6 we obtain an exact sequence
@@
(7.2) $(\Lambda/\mathfrak{f})^{\times}\stackrel{{\scriptstyle\partial}}{{\longrightarrow}}\mathrm{Cl}(\Lambda,\mathfrak{f})\stackrel{{\scriptstyle q_{1}}}{{\longrightarrow}}\mathrm{Cl}(\Lambda)\longrightarrow 0.$
@@ 7.4. Decomposition via central idempotents
Much of the following notation is adopted from [BHJ22]. Let $C=Z(A)$ and let $\mathcal{O}_{C}$ be the integral closure of $\mathcal{O}$ in $C$. Set $\mathfrak{g}:=\mathfrak{f}\cap\mathcal{O}_{C}$ and note that this is a full ideal of $\mathcal{O}_{C}$. Let $e_{1},\ldots,e_{r}$ be the primitive idempotents of $C$ and set $A_{i}=Ae_{i}$. Then
(7.3) $A=A_{1}\oplus\cdots\oplus A_{r}$
is a decomposition of $A$ into indecomposable two-sided ideals (see [CR81, (3.22)]). Each $A_{i}$ is a simple $K$-algebra with identity element $e_{i}$. The centers $K_{i}:=Z(A_{i})$ are finite field extensions of $K$ via $K\rightarrow K_{i}$, $\alpha\mapsto\alpha e_{i}$, and we have $K$-algebra isomorphisms $A_{i}\cong\mathrm{Mat}_{n_{i}}(D_{i})$ where $D_{i}$ is a division algebra with $Z(D_{i})\cong K_{i}$ (see [CR81, (3.28)]). The Wedderburn decomposition (7.3) induces decompositions
$C=K_{1}\oplus\cdots\oplus K_{r}\quad\textrm{and}\quad\mathcal{O}_{C}=\mathcal{O}_{K_{1}}\oplus\cdots\oplus\mathcal{O}_{K_{r}},$
@@
$\mathcal{M}=\mathcal{M}_{1}\oplus\cdots\oplus\mathcal{M}_{r},\quad\mathfrak{f}=\mathfrak{f}_{1}\oplus\cdots\oplus\mathfrak{f}_{r}\quad\textrm{and}\quad\mathfrak{g}=\mathfrak{g}_{1}\oplus\cdots\oplus\mathfrak{g}_{r}.$